Output (units) Total cost (R)

0 100

1 110

2 130

3 166

4 220

5 300

The average variable cost of the third unit is  

 



1
Expert's answer
2021-08-10T11:43:20-0400

TC=FC+VC


"AVC=\frac{VC}{Q}=\frac{66}{3}=22"